# Physical meaning of non-extensive term in Massieu functions

### v2

@@ IV.1 Ideal case
For $N$ non-interacting particles in a volume $V$ at temperature $T$ the ideal canonical partition function is $\mathcal{Z}_{\text{id}}=(V/\lambda^{3})^{N}/N!$, where $\lambda$ is the thermal de Broglie wavelength. The SM Massieu function is
$\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=\ln\mathcal{Z}_{\text{id}}=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N-\frac{1}{2}\ln N$ (40)
where Stirling’s approximation up to order $\ln N$ was used. In the context of thermodynamics, the TH Massieu function is an extensive quantity:
@@
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln N.$ (42)
This result satisfies Eq. (28) (with $c$ neglected) since the mean squared fluctuations of particle number in the ideal system are $\langle(\Delta\hat{N})^{2}\rangle=N$.
@@ IV.2 Non ideal case
In this subsection we numerically check the validity of the results for the non-extensive term of the excess Massieu function when interactions between particles are present.
As before, we analyze the system of particles described by $\hat{U}$, $\hat{N}$ and $V$, where energy and particle number fluctuate and volume is constant. We are interested in the non-extensive term of the Massieu function that corresponds to the excess free energy: $\tilde{S}_{\text{ex}}[\tfrac{1}{T},N]=-\tilde{F}_{\text{ex}}/T$. This quantity is related to the particle fluctuations as can be seen in Eq. (29). The mean squared fluctuations are [see Eq. (24)]
@@
The excess chemical potential is $\mu_{\text{ex}}=\frac{\partial F_{\text{ex}}}{\partial N}$, and the one including non-extensive terms is $\tilde{\mu}_{\text{ex}}=\frac{\partial\tilde{F}_{\text{ex}}}{\partial N}$. Then, taking the derivative of the previous equation, we obtain
@@
where we used $\tilde{\mu}_{\text{ex}}^{\prime}=\mu_{\text{ex}}^{\prime}+O(N^{-2})$, and terms $O(N^{-2})$ are neglected. As mentioned in the introduction, Siepmann et al. Siepmann et al. (1992) analyzed finite size corrections to the excess chemical potential in a particle system with fixed volume and periodic boundary conditions. They found that
@@
Next, we contrast the analytical expression against numerical results for hard disks and spheres. Let us consider first hard disk interaction. Mulero et al. have proposed a simple equation of state (EOS) for a system of hard disks Mulero et al. (2009),
@@
Figure 1: Non extensive term $\frac{\tilde{F}_{\rm ex}}{T}-\frac{F_{\rm ex}}{T}$ against the packing fraction $\eta$ for a system of Hard Disks. The curve corresponds to the rhs of Eq. (45). Symbols represent numerical results; $N$ is the number of particles.
Figure 2: Non extensive term $\frac{\tilde{F}_{\rm ex}}{T}-\frac{F_{\rm ex}}{T}$ against the packing fraction $\eta$ for a system of Hard Spheres. The curve corresponds to Eq. (45). Dots represent numerical results; $N$ is the number of particles.
This average can be numerically obtained via Monte Carlo simulations; see Appendix B. In Fig. 1 we show numerical results for $N=10,15$ and $20$ particles. We can observe that the agreement with the theoretical prediction of Eq. (45) improves as the number of particle increases.
Next we analyse hard sphere interaction. The EOS obtained by Carnahan and Starling Carnahan and Starling (1969) is widely used for hard spheres in the fluid phase. It is given by
@@
In Fig. 2 we show numerical results for $N=10,15$ and $20$ particles. As we observed for hard disks, the agreement with the theoretical line also improves for larger system size.
@@ V Spin lattice
Let us consider a $d$-dimensional lattice with $N$ spins; each spin has two possible states: up or down. There are $\hat{N}_{+}$ and $\hat{N}_{-}=N-\hat{N}_{+}$ spins in the up and down states. The total number of spins, $N$, is constant and plays the same role as the volume in the previous example. Each spin, $s_{i}$, takes the value 1 or $-1$. The microscopic state is given by $\vec{s}=(s_{1},\dots,s_{N})$ and its energy is
$\hat{H}=\hat{E}-h\hat{M}$ (56)
@@
$e^{\tilde{S}(\hat{M},\hat{E})}$ $=\Omega(\hat{M},\hat{E})$ (57)
$e^{\tilde{S}[\hat{M},\frac{1}{T}]}$ $=\sum_{\hat{E}}e^{\tilde{S}(\hat{M},\hat{E})}e^{-\frac{1}{T}\hat{E}}$ (58)
@@ V.1 Ideal case
where terms $O(N^{0})$ are neglected (as in Stirling’s approximation above). The fluctuations are,
@@
where we have used that $\frac{h}{T}=-\frac{\partial S_{\text{id}}(M)}{\partial M}$ (relationship that can be obtained from the Legendre transform $S_{\text{id}}[\frac{h}{T}]=S_{\text{id}}(M)+\frac{h}{T}M$). Therefore, comparing Eqs. (64) and (65) it can be seen that the theory correctly predicts the non-extensive term of the Massieu function at order $\ln N$.
@@ V.2 Non ideal case
Applying the thermodynamic limit approximation to Eq. (59) we have the Legendre transform
@@
$S[M,\tfrac{1}{T}]$ and $S[\tfrac{h}{T},\tfrac{1}{T}]$ are associated to the Helmholtz and Gibbs free energies of the spin lattice, respectively. Let us consider the exact solution of the Ising model of $N$ spins in 1D with periodic boundary conditions; $\ln\mathcal{Z}(\tfrac{h}{T},\tfrac{1}{T})$ is extensive, so $S[\tfrac{h}{T},\tfrac{1}{T}]=\tilde{S}[\tfrac{h}{T},\tfrac{1}{T}]$; see, for example, Sec. 5.6.1 in Reichl (1998) or Sec. 6.5.2 in Schwabl (2006). The result is
@@
where $S_{\text{ex}}[M,\frac{1}{T}]=S[M,\frac{1}{T}]-S_{\text{id}}(M)$, and $S_{\text{id}}(M)$ is given by (63). The mean squared fluctuations are $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$; the ideal fluctuations were already calculated in (66).
@@
In order to evaluate the sum in the last equation we consider a system small enough so that all possible microstates $\omega_{M}$, for a given magnetization $M$, can be computed. The ideal value, $\tilde{S}_{\text{id}}(M)$, is given by Eq. (61) with $\hat{M}=M$. The numerical results are obtained for values of $\epsilon/T$ between 0 and 1.2, and for $M=N/2$. Figure 3 shows results for $N$ equal 16 and 32. As the system size increases, numerical results approach the theoretical prediction of Eq. (V.2). This is an expected behavior since the theoretical results hold for large $N$.
@@ VI Summary and conclusions
In a review on finite system thermodynamics Swendsen (2018), Swendsen suggested that “For future work on the properties of small systems, the most promising approach seems to be a straightforward application of statistical mechanics”. The calculation of non-extensive terms of Massieu functions presented here intends to represent one step in that direction. Non-extensive terms order $\ln N$ are just one ingredient to be taken into account in the analysis of small systems; there are other effects not mentioned here that are analyzed in, for example, Hill (1964); Bedeaux et al. (2020).
The main result is that non-extensive terms are a function of thermodynamic fluctuations. It was obtained from a Taylor expansion around the equilibrium state that holds as long as the system is not close to a phase transition. The result was analytically verified in the ideal case, when interactions can be neglected, and numerically verified in the non-ideal case. The examples analyzed were a particle system with hard sphere interactions and a one-dimensional spin lattice. Numerical simulations were performed on small systems; our result, instead, holds for large $N$. We observed, in both systems analyzed, that the difference between theory and numerical results decreases as $N$ increases, in agreement with our prediction. There is a very good agreement in the close to ideal regime, where interactions are small but are not neglected (corresponding to small concentration $\rho$ in a particle system, or small coupling $\epsilon$ in a spin lattice).
It was shown that the non-extensive term of the excess free energy of a particle system, Eq. (45), is consistent with previous results obtained for the excess chemical potential using the Widom insertion formula, Siepmann et al. (1992); Di Muro and Hoyuelos (2021). There is an important connection between non-extensive terms and transition probabilities that was first explored in Ref. Di Muro and Hoyuelos (2021) in the context of systems that have fluctuations only in the number of particles $N$. This connection will be further explored in future work using the more general results developed here.
@@ Acknowledgements.
Eq. (37) can be used to demonstrate that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant.
We obtain the probability of fluctuations, $\Delta\hat{\vec{X}}=\hat{\vec{X}}-\vec{X}$, directly from the probability of $\hat{\vec{X}}$, Eq. (13). For simplicity, we consider here the ensemble where all extensive variables (except volume) have fluctuations; it is described by the first Massieu function, that is, the entropy $\tilde{S}(\hat{\vec{X}})$. Expanding $\tilde{S}(\hat{\vec{X}})$ in a Taylor series up to second order in $\Delta\hat{\vec{X}}$, we have
@@
therefore, the factor $e^{-\Delta\tilde{S}[\vec{Y}]+\Delta\tilde{S}(\vec{X})}$, a function of the equilibrium parameters $\vec{X}$ and $\vec{Y}$, is the normalization of the multivariate normal distribution given by $(2\pi)^{-r/2}(\det C)^{-1/2}$, where $r$ is the number of variables; see, for example, Mishin (2015) or Sec. 7.C in Reichl (1998); the multivariate Gaussian distribution is a consequence of the central limit theorem, see Ch. V in Khinchin (1949).
@@
where $\textbf{r}_{i}$ is the position of the center of particle $i$. Thus, for a given configuration the exponential in Eq. (52) can be either 0 or 1. The average is then obtained by exploring a large number of particle configurations. A drawback of this method is that for a large number of particles and moderate concentration values, the probability of randomly run into a non-overlapping configuration of particles is vanishingly small. Thus, the number of configurations that need to be explored becomes significantly high.
We employ an alternative approach to tackle this issue. For hard disk or hard sphere interaction the average in Eq. (52) is simply the probability of a non-overlapping configuration when the particles positions are chosen at random. Let us call $P(\bar{I}_{N})$ the probability that none of the $N$ particles overlap or, in other words, that the intersection among them is empty: $P\left(\bigcap_{i=1}^{N}o_{i}=\emptyset\right)$, where $o_{i}$ is the area or volume occupied by particle number $i$. Also, we call $P(\bar{I}_{j}|\bar{I}_{j-1})$ the conditional probability that $j$ particles do not overlap given that $j-1$ do not, or $P\left(\bigcap_{i=1}^{j}o_{i}=\emptyset\big{|}\bigcap_{i=1}^{j-1}o_{i}=\emptyset\right)$. We can write
@@
    Set the positions of the $j-1$ particles in a close packed arrangement, hexagonal for instance.
@@
    Perform a thermalization process, in order to reach a typical equilibrium configuration.
@@
The expressions for fluctuations, external field and Massieu functions for the 1D spin lattice are given in this appendix. The field as a function of $M$ and $1/T$ is
$\frac{h}{T}=\log\left(A+\sqrt{A^{2}+1}\right)$ (78)
@@
And the mean squared fluctuations $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$ are given by

### v3

@@ IV.1 Ideal case
For $N$ noninteracting particles in a volume $V$ at temperature $T$, the ideal canonical partition function is $\mathcal{Z}_{\text{id}}=(V/\lambda^{3})^{N}/N!$, where $\lambda$ is the thermal de Broglie wavelength. The SM Massieu function is
$\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=\ln\mathcal{Z}_{\text{id}}=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N-\frac{1}{2}\ln(2\pi N)$ (41)
where Stirling’s approximation was used. In the context of thermodynamics, the TH Massieu function is an extensive quantity:
@@
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln N-\frac{1}{2}\ln(2\pi).$ (43)
This result satisfies Eq. (29) [with $c=-\frac{1}{2}\ln(2\pi)$], since the mean-squared fluctuations of particle number in the ideal system are $\langle(\Delta\hat{N})^{2}\rangle=N$.
@@ IV.2 Non ideal case
In this section we numerically check the validity of the results for the nonextensive term of the excess Massieu function when interactions between particles are present.
As before, we analyze the system of particles described by $\hat{U}$, $\hat{N}$, and $V$, where energy and particle number fluctuate and volume is constant. We are interested in the nonextensive term of the Massieu function that corresponds to the excess free energy: $\tilde{S}_{\text{ex}}[\tfrac{1}{T},N]=-\tilde{F}_{\text{ex}}/T$. This quantity is related to the particle fluctuations, as can be seen in Eq. (30). The mean-squared fluctuations are [see Eq. (25)]
@@
The excess chemical potential is $\mu_{\text{ex}}=\frac{\partial F_{\text{ex}}}{\partial N}$, and the one including nonextensive terms is $\tilde{\mu}_{\text{ex}}=\frac{\partial\tilde{F}_{\text{ex}}}{\partial N}$. Then, taking the derivative of the previous equation, we obtain
@@
where we used $\tilde{\mu}_{\text{ex}}^{\prime}=\mu_{\text{ex}}^{\prime}+O(N^{-2})$, and terms $O(N^{-2})$ are neglected. As mentioned in the Introduction, Siepmann et al. Siepmann et al. (1992) analyzed finite-size corrections to the excess chemical potential in a particle system with fixed volume and periodic boundary conditions. They found that
@@
Next, we contrast the analytical expression against numerical results for hard disks and spheres. Let us consider first hard disk interaction. Solana have proposed a simple equation of state (EOS) for a system of hard disks Mulero et al. (2009),
@@
Figure 1: Nonextensive term $\tilde{F}_{\rm ex}/T$ (dots) and $F_{\rm ex}/T$ (solid line) as a function of the packing fraction $\eta$ for a system of hard disks. The dashed lines represent $F_{\rm ex}/T-1/2\log{\Gamma}$ [see Eq. (46)]. $N$ is the number of particles.
Figure 2: Nonextensive term $\tilde{F}_{\rm ex}/T-F_{\rm ex}/T$ against the packing fraction $\eta$ for a system of hard disks. The curve corresponds to the right hand side of Eq. (46). Symbols represent numerical results; $N$ is the number of particles. The statistical error does not exceed $0.01\%$.
Figure 3: Nonextensive term $\tilde{F}_{\rm ex}/T$ (dots) and $F_{\rm ex}/T$ (solid line) against the packing fraction $\eta$ for a system of hard spheres. The dashed lines represent $F_{\rm ex}/T-1/2\log{\Gamma}$ [see Eq. (46)]. $N$ is the number of particles.
Figure 4: Nonextensive term $\tilde{F}_{\rm ex}/T-F_{\rm ex}/T$ against the packing fraction $\eta$ for a system of hard spheres. The curve corresponds to Eq. (46). Dots represent numerical results; $N$ is the number of particles. The statistical error is smaller than $0.01\%$.
This average can be numerically obtained via Monte Carlo simulations; see Appendix B.
In Fig. 1 we compare the thermodynamic and statistical mechanics excess free energy, $\tilde{F}_{\text{ex}}$ and $F_{\text{ex}}$, respectively, for $N=10,15$, and $20$ particles The dashed curve is the correction to $F_{\text{ex}}$, given by Eq. (46). In Fig. 2 we show the correction given by Eq. (46). We can observe in Fig. 2 that the agreement with the theoretical prediction of Eq. (46) improves as the number of particles increases. For instance, for $\eta=0.4$ the deviations are around $8\%$, $5\%$, and $4\%$ for $N=10$, $15$, and $20$, respectively.
Next we analyze hard sphere interaction. The EOS obtained by Carnahan and Starling Carnahan and Starling (1969) is widely used for hard spheres in the fluid phase. It is given by
@@
In Fig. 3 we show numerical results for the SM and TH free energy, for $N=10,15$, and $20$ particles. In Fig. 4 we can see in more detail the correction of Eq. (46). As we observed before for hard disks, the agreement with the theoretical line also improves for larger system size. Here, for $\eta=0.2$ the deviations are around $7\%$, $5\%$, and $4\%$ for $N=10$, $15$, and $20$, respectively. Other EOS were also considered Boublik (1986); Kolafa et al. (2004), but no significant difference was found.
@@ V Spin lattice
Let us consider a $d$-dimensional lattice with $N$ spins, where each spin has two possible states: up or down. There are $\hat{N}_{+}$ and $\hat{N}_{-}=N-\hat{N}_{+}$ spins in the up and down states. The total number of spins, $N$, is constant and plays the same role as the volume in the previous example. Each spin, $s_{i}$, takes the value 1 or $-1$. The microscopic state is given by $\vec{s}=(s_{1},\dots,s_{N})$ and its energy is
$\hat{H}=\hat{E}-h\hat{M},$ (57)
@@
$e^{\tilde{S}(\hat{M},\hat{E})}$ $=\Omega(\hat{M},\hat{E}),$ (58)
$e^{\tilde{S}[\hat{M},\frac{1}{T}]}$ $=\sum_{\hat{E}}e^{\tilde{S}(\hat{M},\hat{E})}e^{-\frac{1}{T}\hat{E}},$ (59)
@@ V.1 Ideal case
where the terms $O(N^{0})$ are neglected (as in Stirling’s approximation above). The fluctuations are
@@
where we have used that $\frac{h}{T}=-\frac{\partial S_{\text{id}}(M)}{\partial M}$ (a relationship that can be obtained from the Legendre transform $S_{\text{id}}[\frac{h}{T}]=S_{\text{id}}(M)+\frac{h}{T}M$). Therefore, comparing Eqs. (65) and (66), it can be seen that the theory correctly predicts the nonextensive term of the Massieu function at order $\ln N$.
@@ V.2 Nonideal case
By applying the thermodynamic limit approximation to Eq. (60), we have the Legendre transform
@@
$S[M,\tfrac{1}{T}]$ and $S[\tfrac{h}{T},\tfrac{1}{T}]$ are associated to the Helmholtz and Gibbs free energies of the spin lattice, respectively. Let us consider the exact solution of the Ising model of $N$ spins in one dimension with periodic boundary conditions; $\ln\mathcal{Z}(\tfrac{h}{T},\tfrac{1}{T})$ is extensive, so $S[\tfrac{h}{T},\tfrac{1}{T}]=\tilde{S}[\tfrac{h}{T},\tfrac{1}{T}]$ (see, for example, Sec. 5.6.1 in Reichl (1998) or Sec. 6.5.2 in Schwabl (2006)). The result is
@@
where $S_{\text{ex}}[M,\frac{1}{T}]=S[M,\frac{1}{T}]-S_{\text{id}}(M)$, and $S_{\text{id}}(M)$ is given by (64). The mean-squared fluctuations are $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$; the ideal fluctuations were already calculated in (67).
@@
In order to evaluate the sum in the last equation, we consider a system small enough so that all possible microstates $\omega_{M}$, for a given magnetization $M$, can be computed. The ideal value, $\tilde{S}_{\text{id}}(M)$, is given by Eq. (62) with $\hat{M}=M$. The numerical results are obtained for values of $\epsilon/T$ between 0 and 1.2, and for $M=N/2$. Figures 5 and 6 show results for $N$ equal 16 and 32. Figure 5 shows $S_{\text{ex}}[M,\tfrac{1}{T}]$ and $\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$ as functions of $\epsilon/T$ with continuous and dashed curves, respectively; the dots represent numerical results for the SM Massieu function $\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$. The significance of the nonextensive term is given by the difference between these curves; the percentage of this correction, for $\epsilon/T=1$, is about 10% for $N=16$ and 5% for $N=32$. Figure 6 shows with more detail the accuracy of the theoretical prediction for the difference $\Delta\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$, see Eq. (V.2). Theoretically, this difference is an intensive quantity independent of the system size $N$, while the numerical results depend on $N$, showing a better agreement with the theory for larger $N$. This is an expected behavior since the theoretical results hold for large $N$.
Figure 5: TH Massieu function $S_{\text{ex}}[M,\tfrac{1}{T}]$ [continuous curves) and SM Massieu function $\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$ (dashed curves, using the correction of Eq. (V.2)] against $\epsilon/T$ for the one-dimensional spin lattice. Blue curves are for $N=16$, green curves, for $N=32$. Dots correspond to the numerical evaluation of $\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$; in both cases, an external field is applied such that $M=N/2$.
@@ VI Summary and conclusions
In a review on finite system thermodynamics Swendsen (2018), Swendsen suggested that “For future work on the properties of small systems, the most promising approach seems to be a straightforward application of statistical mechanics.” The calculation of nonextensive terms of Massieu functions presented here intends to represent one step in that direction. nonextensive terms order $\ln N$ are just one ingredient to be taken into account in the analysis of small systems; there are other effects not mentioned here that are analyzed in, for example, Hill (1964); Bedeaux et al. (2020).
The nonextensive term of any Massieu function can be written in terms of the determinant of the fluctuation matrix; see Eq. (38). The result was analytically verified in the ideal case, when interactions can be neglected, and numerically verified in the nonideal case. The examples analyzed were particle systems with hard disk and hard sphere interactions and a one-dimensional spin lattice. The procedure to obtain the nonextensive term is based on a Taylor expansion of equilibrium fluctuations where terms of order larger than 2 are neglected; this approximation is justified for large $N$, since the $i$th derivative of a (extensive) Massieu function respect to $i$ extensive variables is $O(N^{-i+1})$ and the central moments are $O(N)$; see Eq. (23), or Eq. (III.2) for the general derivation. For this reason we expect a better performance of the results for larger values of $N$ and large departures close to a critical point where fluctuations diverge. We observed that the difference between theory and numerical results decreases as $N$ increases, in agreement with this prediction. An increase in $N$ also implies, in our examples, an increase in the system size that reduces the implicit or anomalous finite-size effects caused by the periodic boundary conditions.
Numerical simulations show that there is a regime, characterized by weak interactions (corresponding to small concentration $\rho$ in a particle system, or small coupling $\epsilon$ in a spin lattice), for which the prediction for the nonextensive term holds even for quite small values of $N$. Discrepancies with the theory appear for large concentration in particle systems or for strong interaction in the spin lattice. A criterion to determine whether a system is small or large is to compare the system size with the fluctuation correlation length. Implicit finite-size corrections become relevant if the correlation length is similar to the system size; we consider that this is the cause of the discrepancies that appear for stronger interactions (large densities for the particle systems or large coupling for the spin lattice). This argumentation, based on the correlation length criterion presented in Lebowitz and Percus (1961), is qualitative; further research is needed to quantitatively determine the validity range of the results.
It was shown that the nonextensive term of the excess free energy of a particle system, Eq. (46), is consistent with previous results obtained for the excess chemical potential using the Widom insertion formula Siepmann et al. (1992); Di Muro and Hoyuelos (2021). There is an important connection between nonextensive terms and transition probabilities that was first explored in Ref. Di Muro and Hoyuelos (2021) in the context of systems that have fluctuations only in the number of particles $N$. This connection will be further explored in future work using the more general results developed here.
@@ Acknowledgements.
Equation (38) can be used to demonstrate that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant.
We obtain the probability of fluctuations, $\Delta\hat{\vec{X}}=\hat{\vec{X}}-\vec{X}$, directly from the probability of $\hat{\vec{X}}$, Eq. (14). For simplicity, we consider here the ensemble where all extensive variables (except volume) have fluctuations ($s=r$); it is described by the first Massieu function, that is, the entropy $\tilde{S}(\hat{\vec{X}})$. Expanding $\tilde{S}(\hat{\vec{X}})$ in a Taylor series up to second order in $\Delta\hat{\vec{X}}$, we have
@@
and therefore, the factor $e^{-\Delta\tilde{S}[\vec{Y}]+\Delta\tilde{S}(\vec{X})}$, a function of the equilibrium parameters $\vec{X}$ and $\vec{Y}$, is the normalization of the multivariate normal distribution given by $(2\pi)^{-r/2}(\det C)^{-1/2}$, where $r$ is the number of variables. See, for example, Mishin (2015) or Sec. 7.C in Reichl (1998); the multivariate Gaussian distribution is a consequence of the central limit theorem, see Ch. V in Khinchin (1949).
@@
where $\textbf{r}_{i}$ is the position of the center of particle $i$. Thus, for a given configuration the exponential in Eq. (53) can be either 0 or 1. The average is then obtained by exploring a large number of particle configurations. A drawback of this method is that for a large number of particles and moderate concentration values, the probability of randomly obtaining a non-overlapping configuration of particles is vanishingly small. Thus, the number of configurations that need to be explored becomes significantly high.
We employ an alternative approach to tackle this issue. For a hard disk or hard sphere interaction the average in Eq. (53) is simply the probability of a non-overlapping configuration when the particle positions are chosen at random. Let us call $P(\bar{I}_{N})$ the probability that none of the $N$ particles overlap or, in other words, that the intersection among them is empty: $P\left(\bigcap_{i=1}^{N}o_{i}=\emptyset\right)$, where $o_{i}$ is the area or volume occupied by particle number $i$. Also, we call $P(\bar{I}_{j}|\bar{I}_{j-1})$ the conditional probability that $j$ particles do not overlap given that $j-1$ do not, or $P\left(\bigcap_{i=1}^{j}o_{i}=\emptyset\big{|}\bigcap_{i=1}^{j-1}o_{i}=\emptyset\right)$. We can write
@@
    Set the positions of the $j-1$ particles in a close packed arrangement, hexagonal, for instance.
@@
    Perform a thermalization process in order to reach a typical equilibrium configuration.
@@
The expressions for fluctuations, external field and Massieu functions for the one-dimensional spin lattice are given in this appendix. The field as a function of $M$ and $1/T$ is
$\frac{h}{T}=\log\left(A+\sqrt{A^{2}+1}\right),$ (79)
@@
And the mean-squared fluctuations $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$ are given by